### v2

# Vireo: Leveraging Depth and Language for Open-Vocabulary Domain-Generalized Semantic Segmentation

Abstract
Open-Vocabulary semantic segmentation (OVSS) and domain generalization in semantic segmentation (DGSS) highlight a subtle complementarity that motivates Open-Vocabulary Domain-Generalized Semantic Segmentation (OV-DGSS). OV-DGSS aims to generate pixel-level masks for unseen categories while maintaining robustness across unseen domains, a critical capability for real-world scenarios such as autonomous driving in adverse conditions. We introduce Vireo, a novel single-stage framework for OV-DGSS that unifies the strengths of OVSS and DGSS for the first time. Vireo builds upon the frozen Visual Foundation Models (VFMs) and incorporates scene geometry via Depth VFMs to extract domain-invariant structural features. To bridge the gap between visual and textual modalities under domain shift, we propose three key components: (1) GeoText Prompts, which align geometric features with language cues and progressively refine VFM encoder representations; (2) Coarse Mask Prior Embedding (CMPE) for enhancing gradient flow for faster convergence and stronger textual influence; and (3) the Domain-Open-Vocabulary Vector Embedding Head (DOV-VEH), which fuses refined structural and semantic features for robust prediction. Comprehensive evaluation on these components demonstrates the effectiveness of our designs. Our proposed Vireo achieves the state-of-the-art performance and surpasses existing methods by a large margin in both domain generalization and open-vocabulary recognition, offering a unified and scalable solution for robust visual understanding in diverse and dynamic environments. Code is available at https://github.com/anonymouse-9c53tp182bvz/Vireo.
1 Introduction
Open-Vocabulary Domain-Generalized Semantic Segmentation (OV-DGSS) denotes the joint execution of open-vocabulary semantic segmentation (OVSS)[1, 2, 3, 4] and domain generalization in semantic segmentation (DGSS)[5, 6, 7, 8] tasks. It involves training a model that, without access to target-domain samples or annotations for novel categories, can generate pixel-wise segmentation for unseen classes while sustaining high performance across previously unseen domains (such as different cities, lighting environments, or climatic conditions). For instance, when an autonomous vehicle receives a query like “Can I park next to that bollard?”, its perception system must comprehend the linguistic input and accurately segment the referenced object at the pixel level—even under adverse conditions such as dim lighting, rain-streaked lenses, or region-specific visual appearances.
OVSS and DGSS share notable similarities and can both be implemented using either multi-stage or single-stage strategies. In multi-stage OVSS [1], candidate regions or coarse masks are generated and then classified by a text approach, while multi-stage DGSS [9]first aligns domains through adversarial alignment or style transfer and then trains the segmentation approach on the aligned features.
In single-stage OVSS [10, 11], the segmentation head is dynamically conditioned on text prompts to directly produce masks for each class. In contrast, single-stage DGSS [12, 13] incorporates domain-invariant modules within the backbone or segmentation head, enabling the model to jointly learn segmentation and generalization in a unified forward pass without requiring distinct stages. The key distinction lies in their focus: OVSS requires integration of visual features with textual semantics to accurately identify unseen classes, whereas DGSS emphasizes robustness to domain shifts.
Consequently, integrating open-vocabulary recognition of novel classes with domain robustness into a unified framework presents two primary challenges: (1) Text–vision alignment modules often degrade outside the source domain, leading to significant performance drops even for previously seen classes. (2) Domain-invariant strategies may suppress fine-grained semantic cues, hindering the model’s ability to precisely respond to detailed textual queries.
Figure 1: Overview of Vireo and effectiveness. (I) Performance comparison under various adverse conditions across ACDC and DELIVER datasets, showing that Vireo consistently outperforms existing methods. (II) Qualitative visualization of segmentation predictions (left) and attention maps (right) under extreme night scenes, illustrating Vireo’s robustness and precise alignment with semantic cues. (III) Architectural comparison: (a) Traditional OVSS and (b) DGSS pipelines freeze or fine-tune VFM separately without cross-modal integration; (c) Our proposed Vireo introduces GeoText Prompts and Depth-VFM integration to enhance both semantic alignment and domain robustness.
Recent single-stage DGSS studies have increasingly adopted strategies that fine-tune learnable tokens across the layers of a frozen Visual Foundation Model (VFM) [14, 15, 16, 17, 18] to adapt its feature representations. In OVSS, the VFM encoder is typically fully frozen, with efforts focused on designing the decoder to endow the model with open-vocabulary recognition capabilities. This reveals a subtle complementarity between the two paradigms: DGSS emphasizes the encoder by leveraging the VFM’s strong feature generalization to learn cross-domain representations, whereas OVSS freezes the encoder and emphasizes the decoder to enable open-vocabulary recognition.
Moreover, in cross-domain scenarios, depth and geometric cues are largely insensitive to variations in illumination and texture[19, 20]. They supply reliable spatial constraints, easing the distribution shift of RGB features and sharpening boundary localization. Recent studies such as DepthForge[21] have shown that injecting depth prompts into a frozen VFM boosts domain generalization. Motivated by these findings, we adopt DepthAnything V2 as our Depth VFM: its diverse pre-training delivers consistent depth estimation across domains, and keeping it fully frozen incurs minimal training cost while preserving real-time inference speed.
In this paper, we propose a VFM-based single-stage pipeline for OV-DGSS, termed Vireo. Specifically, in the encoder, both the VFM and DepthAnything modules are kept frozen. The VFM is leveraged to robustly capture cross-domain visual features, while DepthAnything extracts the scene’s intrinsic geometric structure. On this basis, we introduce Geometric-Text Prompts (GeoText Prompts) to fuse the extracted structural features with manually provided textual cues, progressively refining the feature maps between the frozen VFM layers. To mitigate the slow convergence caused by sparse gradients propagating through the frozen encoder, we introduce a Coarse Mask Prior Embedding (CMPE) at the beginning of the decoder to inject denser gradient signals. This design accelerates the convergence of mask supervision and further strengthens the influence of textual priors. Subsequently, we design the Domain-Open-Vocabulary Vector Embedding Head (DOV-VEH) to strengthen the synergistic integration of structural and textual modalities, ensuring that both cross-domain structural features and open-vocabulary cues learned by the GeoText prompts are fully utilized in the final prediction.
For challenge 1: We find that GeoText Prompts not only capture structural and semantic cues within the frozen VFM encoder but also guide the progressive refinement of its feature representations. For challenge 2: The CMPE enhances gradient back-propagation into the encoder, while the redesigned DOV-VEH deepens the fusion of visual and textual priors. These components form a unified framework that simultaneously achieves domain robustness and strong open-vocabulary generalization. Our main contributions are summarized as follows:
(1) We propose Vireo, a novel VFM-based single-stage framework for OV-DGSS.
(2) We introduce GeoText Prompts to progressively refine frozen VFM features by injecting geometric cues from DepthAnything and aligning them with textual semantics, enabling structural-semantic fusion across encoder layers.
(3) We design two complementary modules—CMPE for enhancing gradient flow and DOV-VEH for fusing visual and textual priors—together boosting segmentation performance under domain shifts and unseen classes.
Figure 2: Overview of the proposed Vireo framework for OV-DGSS. (a) Framework Overview. Geometric-Text Prompts permeate our model: they’re injected in Tunable Vireo to align domain priors, reused in CMPE to guide multi-scale feature fusion, and employed in DOV-VEH as queries for open-vocabulary segmentation, forming a unified end-to-end loop. (b) Tunable Vireo. GeoText Prompts are used to inject structural-textual priors across multiple layers. Depth-aware fusion and attention optimization are applied between intermediate frozen layers to progressively refine representations. (c) Coarse Mask Prior Embedding (CMPE). Multi-scale features from VFM are combined with GeoText Prompts to provide dense supervision and gradient signals for downstream modules. (d) Domain-Open-Vocabulary Head (DOV-VEH). Multi-level features are processed by a Pixel Decoder and Transformer Decoder to produce final predictions guided by open-vocabulary text queries.
2 Related Works
Open-Vocabulary Semantic Segmentation. Open-Vocabulary Semantic Segmentation (OVSS) aims to segment objects based on arbitrary textual descriptions, moving beyond fixed, pre-defined categories. A key enabler for OVSS [22, 23, 1, 24] has been vision-language models (VLMs), particularly CLIP [16], which aligns visual and textual representations.
Domain-Generalized Semantic Segmentation. Domain Generalization in Semantic Segmentation (DGSS) addresses the performance degradation when models encounter unseen target domains due to domain shift—variations in data distributions (e.g., lighting, weather). Data augmentation [25, 26] and learning domain-invariant representations [27, 28] are two key strategies to enhance model robustness. This detailed discussion in the appendix further contextualizes our specific contributions within the broader landscape of these research areas.
3 Methodology - Vireo
3.1 Overview
As illustrated in Figure 2, our framework consists of three key modules tailored for the OV-DGSS:
  • Tunable Vireo with GeoText Prompts: Introduces GeoText Prompts to inject and refine both geometric and textual information across layers of a frozen VFM.
  • Coarse Mask Prior Embedding (CMPE): Generates coarse prior masks to guide segmentation and reinforce gradient flow from the decoder back to the frozen encoder.
  • Domain-Open-Vocabulary Vector Embedding Head (DOV-VEH): Integrates visual, geometric, and semantic features to produce final OV-DGSS predictions.
The input RGB image $I\in\mathbb{R}^{H\times W\times 3}$ is duplicated and independently fed into two frozen encoders: a VFM encoder $\mathcal{F}^{V}$ and a DepthAnything encoder $\mathcal{F}^{D}$. These backbones extract a series of multi-scale features $f^{V}_{l}$ and $f^{D}_{l}$ respectively, where $l$ denotes the encoder layer index in $[1,L]$. In parallel, a set of class labels $\mathcal{C}=\{c_{1},\dots,c_{K}\}$ is transformed into natural language prompts $T=\{\texttt{prompt}_{1},\dots,\texttt{prompt}_{K}\}$, which are encoded using a frozen CLIP text encoder $\mathcal{F}^{T}$ to obtain textual embeddings $t_{k}\in\mathbb{R}^{d}$. These features serve as shared semantic priors across the entire framework and are precomputed once at initialization. In the Tunable Vireo module, each layer receives a tuple $(f^{V}_{l},f^{D}_{l},t_{k})$, and applies the GeoText Prompts $P_{l}$ to align and refine the visual features. This is achieved by computing cross-modal attention maps $\mathcal{A}_{l}=\text{Attn}(P_{l},f^{V}_{l},f^{D}_{l},t_{k})$ followed by feature fusion and projection. The refined output $\hat{f^{V}}_{l}$ is then forwarded back to update the VFM’s layer-wise activations.
We select the refined visual features $\{{f}^{V}_{l_{i}}\}_{i=1}^{4}$ from the VFM encoder at layers $l_{1}=8$, $l_{2}=12$, $l_{3}=16$, and $l_{4}=24$ for downstream decoding. In the Coarse Mask Prior Embedding (CMPE) module, these features are first upsampled to a unified spatial resolution and passed through a channel-spatial attention gating function $\mathcal{G}(\cdot)$. The gated outputs are then fused to form a global coarse feature $f^{M}\in\mathbb{R}^{d}$. The $f^{M}$ is projected and matched with the text embeddings $t_{k}\in\mathbb{R}^{d}$ to produce a coarse class probability map $\mathcal{M}\in\mathbb{R}^{H\times W\times K}$, where $\mathcal{M}$ serves both as a weak supervision signal and as a prior to construct the query embeddings $\mathcal{Q}$. The query prior is added to the GeoText Prompts and forwarded to the final segmentation head. In the Domain-Open-Vocabulary Vector Embedding Head (DOV-VEH), the multi-scale features $\{\hat{f}^{V}_{l_{i}}\}_{i=1}^{4}$ are passed through a pixel decoder $\mathcal{D}_{p}(\cdot)$ to enhance spatial representation, followed by a Transformer decoder $\mathcal{D}_{T}(\cdot)$ that leverages positional embedding. The GeoText Prompts serve as learnable queries and interact with both the decoded features and the text embeddings, producing pixel-level mask embeddings $\mathcal{E}_{\text{mask}}(x,y)\in\mathbb{R}^{d}$ and classification embeddings $\mathcal{E}_{\text{cls}}(k)\in\mathbb{R}^{d}$. The final prediction is ^$\hat{\mathcal{M}}(x,y,k)\in\mathbb{R}^{H\times W\times K}$ provides the pixel-wise semantic prediction with both fine-grained detail and open-vocabulary generalization capability.
3.2 Tunable Vireo with GeoText Prompts
To improve efficiency, we first precompute and share the textual prompt embeddings using a frozen CLIP text encoder $\mathcal{F}^{T}$. Specifically, a set of class labels $\mathcal{C}=\{c_{1},\dots,c_{K}\}$ is transformed into language prompts $T=\{\texttt{prompt}_{1},\dots,\texttt{prompt}_{K}\}$, which are encoded as $t_{k}=\mathcal{F}^{T}(\texttt{prompt}_{k})$, where $t_{k}\in\mathbb{R}^{d}$. These embeddings are reused across all GeoText Prompt layers, CMPE, and DOV-VEH modules to avoid redundant computation.
During inference, the input image $I\in\mathbb{R}^{H\times W\times 3}$ is simultaneously processed by a frozen visual encoder $\mathcal{F}^{V}$ and a frozen depth encoder $\mathcal{F}^{D}$ (e.g., DepthAnything). For each selected layer $l\in\{1,\dots,L\}$, we obtain the visual feature map $f^{V}_{l}=\mathcal{F}^{V}_{l}(I)$ and the depth feature map $f^{D}_{l}=\mathcal{F}^{D}_{l}(I)$. Each layer of the Tunable Vireo module receives the tuple $(f^{V}_{l},f^{D}_{l},\{t_{k}\})$ along with a layer-specific GeoText Prompt $P_{l}\in\mathbb{R}^{N\times d}$, where $N$ is the number of learnable prompts.
The prompt $P_{l}$ first interacts with the textual embeddings via a fusion block, then attends to both $f^{V}_{l}$ and $f^{D}_{l}$ through cross-attention mechanisms:
$\mathcal{A}_{l}=\text{CrossAttn}(P_{l},\ f^{V}_{l},\ f^{D}_{l},\ \{t_{k}\}).$ (1)
The attention outputs are fused via weighted summation, then passed through an MLP projection layer and multiplied element-wise with $P_{l}$. A residual connection adds this result to the original feature map $f^{V}_{l}$, yielding a refined visual representation $\hat{f}^{V}_{l}$. Finally, another MLP transforms $\hat{f}^{V}_{l}$ into the input for the next VFM layer, and the updated GeoText Prompt $P_{l+1}$ is passed forward.
This progressive refinement continues across all selected layers, enabling the model to inject and align geometric and semantic information at multiple scales, thereby enhancing cross-domain robustness and open-vocabulary generalization.
3.3 Coarse Mask Prior Embedding (CMPE)
We select the refined visual features $\{\hat{f}^{V}_{l_{i}}\}_{i=1}^{4}$ from the VFM encoder at layers $l_{1}=8$, $l_{2}=12$, $l_{3}=16$, and $l_{4}=24$, respectively. Each feature map is upsampled to a common spatial resolution $(H\times W)$ via bilinear interpolation, and then passed through an Adaptive Attention Gate (AAG) $\mathcal{G}(\cdot)$, which enhances informative channels and spatial regions. Specifically, AAG applies two $1\times 1$ convolutions followed by ReLU and Sigmoid for channel attention, and a $3\times 3$ convolution followed by Sigmoid for spatial attention.
The attended features are concatenated along the channel axis and fused via a $1\times 1$ convolution to restore the embedding dimension $d$, yielding a fused feature representation: $f^{M}=\text{Fuse}(\mathcal{G}(\hat{f}^{V}_{l_{i}}))$. We apply a residual addition between $f^{M}$ and the final layer output $\hat{f}^{V}_{l_{4}}$ to obtain the updated mask feature: $f^{M}=f^{M}+\hat{f}^{V}_{l_{4}}$. This fused feature $f^{M}(x,y)\in\mathbb{R}^{d}$ is projected to the same dimension as the text embeddings $t_{k}\in\mathbb{R}^{d}$ and compared via Einstein summation to generate a coarse semantic probability map $\mathcal{M}(x,y,k)=\left\langle f^{M}(x,y),\ t_{k}\right\rangle$ , where $\mathcal{M}\in\mathbb{R}^{B\times K\times H\times W}$. This coarse mask is supervised with a segmentation loss to enhance gradient flow through the frozen encoder.
To generate query priors for the downstream segmentation head, we first normalize $\mathcal{M}$ across the spatial domain to derive attention weights:
$\alpha_{k}(x,y)=\frac{\exp(\mathcal{M}(x,y,k))}{\sum_{x^{\prime},y^{\prime}}\exp(\mathcal{M}(x^{\prime},y^{\prime},k))}.$ (2)
Then, we compute the class-specific aggregated feature by spatially weighting $f^{M}$:
$f^{\text{class}}_{k}=\sum_{x,y}\alpha_{k}(x,y)\cdot f^{M}(x,y).$ (3)
Each $f^{\text{class}}_{k}$ is projected into the embedding space as $e^{\text{class}}_{k}\in\mathbb{R}^{d}$, and combined with a set of learnable query vectors $\{q_{j}\}_{j=1}^{N_{q}}$ to produce the query priors:
$q^{\text{prior}}_{j}=\sum_{k=1}^{K}\text{Softmax}(\langle q_{j},\ e^{\text{class}}_{k}\rangle)\cdot e^{\text{class}}_{k}.$ (4)
The final query priors $\{q^{\text{prior}}_{j}\}_{j=1}^{N_{q}}$ are added to the corresponding GeoText Prompts and forwarded into the Domain-Open-Vocabulary Vector Embedding Head (DOV-VEH).
3.4 Domain-Open-Vocabulary Vector Embedding Head (DOV-VEH)
The DOV-VEH module receives the refined multi-scale features $\{\hat{f}^{V}_{l_{i}}\}_{i=1}^{4}$ from the VFM encoder, where $l_{1}=8$, $l_{2}=12$, $l_{3}=16$, and $l_{4}=24$, along with the updated GeoText Prompts $\{P_{l}\}$. These features are first processed by a pixel decoder $\mathcal{D}_{p}(\cdot)$, which leverages multi-scale cross-attention to extract rich spatial context: $f^{\text{pix}}=\mathcal{D}_{p}\left(\hat{f}^{V}_{l_{i}}\right)$. The fused feature $f^{\text{pix}}\in\mathbb{R}^{H\times W\times d}$ is then compressed via a $1\times 1$ convolution and enriched with sinusoidal positional encoding to preserve spatial structure.
The enhanced features are fed into a Transformer Decoder $\mathcal{D}_{T}(\cdot)$, where the GeoText Prompts act as learnable queries. Through stacked layers of self-attention and cross-attention with $f^{\text{pix}}$, the model captures fine-grained visual-semantic alignment, yielding a set of high-resolution mask features $\mathcal{E}_{\text{mask}}(x,y)\in\mathbb{R}^{d}$ at each spatial position.
Simultaneously, the GeoText Prompts $\{P\}$ are passed through a two-layer MLP and then interact with the text embeddings $\{t_{k}\}$ (precomputed from the CLIP text encoder) to produce classification-level representations $\mathcal{E}_{\text{cls}}(k)\in\mathbb{R}^{d}$. The final segmentation prediction ^$\hat{\mathcal{M}}\in\mathbb{R}^{H\times W\times K}$ is generated via an Einstein summation over the two embeddings:
^$\hat{\mathcal{M}}(x,y,k)=\sum_{d=1}^{D}\mathcal{E}_{\text{mask}}(x,y,d)\cdot\mathcal{E}_{\text{cls}}(k,d),$ (5)
where $D$ is the feature embedding dimension. This design enables DOV-VEH to generate pixel-level segmentation masks that are both spatially accurate and semantically aligned with open-vocabulary textual queries.
Table 1: Performance comparison between our Vireo and existing OVSS and DGSS methods under Citys. $\rightarrow$ ACDC + BDD. + Map., and GTA5. $\rightarrow$ Citys. + BDD. + Map. generalization settings. Top three results are highlighted as best, second, and third, respectively. (%)
Method Proc. & Year Trained on Cityscapes Trained on GTA5
Night-ACDC Fog-ACDC Rain-ACDC Snow-ACDC BDD 100k Mapillary GTA5 Cityscapes BDD100k Mapillary
OVSS Method
FC-CLIP [29] NeurIPS2023 40.8 64.4 63.2 61.5 55.92 66.12 47.12 53.54 51.41 58.60
EBSeg [30] CVPR2024 27.7 56.5 51.8 50.1 48.91 63.40 42.61 44.80 40.59 56.28
CAT-Seg [10] CVPR2024 37.2 58.3 45.6 49.0 48.26 54.74 45.18 43.52 44.28 50.88
SED [3] CVPR2024 38.7 69.0 56.4 60.2 53.30 64.32 48.93 47.45 48.16 57.38
DGSS Method
ResNet based:
IBN [31] ECCV2018 21.2 63.8 50.4 49.6 48.56 57.04 45.06 - - -
RobustNet [9] CVPR2021 24.3 64.3 56.0 49.8 50.73 58.64 45.00 36.58 35.20 40.33
WildNet [32] CVPR2022 12.7 41.2 34.2 28.4 50.94 58.79 47.01 44.62 38.42 46.09
Transformer based:
HGFormer [33] CVPR2023 52.7 69.9 72.0 68.6 53.40 66.90 51.30 - - -
CMFormer [34] AAAI2024 33.7 77.8 67.6 64.3 59.27 71.10 58.11 55.31 49.91 60.09
VFM based:
REIN [35] CVPR2024 55.9 79.5 72.5 70.6 63.54 74.03 62.41 66.40 60.40 66.10
FADA [36] NeurIPS2024 57.4 80.2 75.0 73.5 65.12 75.86 63.78 68.23 61.94 68.09
OV-DGSS Method
Vireo (Ours) - 60.6 82.3 76.3 76.2 66.73 75.99 67.86 70.69 62.91 69.63
4 Experiments
4.1 Datasets & Evaluation Protocols
We evaluate Vireo on six real-world datasets (Cityscapes [37], BDD100K [38], Mapillary [39], ACDC [40], ADE150 [41], and ADE847 [41]) and two synthetic datasets (GTA5 [42] and DELIVER [43]). Cityscapes (City.) is an autonomous-driving dataset with 2,975 training images and 500 validation images, each at a resolution of $2048\times 1024$. BDD100K (BDD.) and Mapillary (Map.) provide 1,000 and 2,000 validation images, respectively, at resolutions of $1280\times 720$ and $1920\times 1080$. ACDC offers 406 validation images captured under extreme conditions (night, snow, fog, and rain), each at $1920\times 1080$. GTA5 is a synthetic dataset that contains 24,966 labeled images obtained from a video game. DELIVER is a multimodal synthetic dataset comprising 3,983 training images, 2,005 validation images, and 1,897 test images across five weather conditions (cloudy, foggy, night, rainy, and sunny); every image is $1042\times 1042$ and there are 25 classes. ADE150 and ADE847 refer to subsets of the ADE20K dataset [41], each containing 2,000 validation images of variable resolution sourced from diverse scenes such as SUN and Places, covering 150 and 847 semantic categories, respectively.
Following the existing DGSS evaluation protocol, we train on one dataset as the source domain and validate on multiple unseen target domains. The three standard evaluation setups are: (1) Cityscapes $\rightarrow$ ACDC; (2) GTA5 $\rightarrow$ Cityscapes, BDD100K, Mapillary; (3) Cityscapes $\rightarrow$ BDD100K, Mapillary, GTA5. To assess our proposed OV-DGSS approaches and compare its open-vocabulary capability against OVSS approaches, we additionally introduce two more configurations: (4) Cityscapes $\rightarrow$ DELIVER, ADE150, ADE847; (5) GTA5 $\rightarrow$ DELIVER, ADE150, ADE847. The evaluation metric is mean Intersection over Union (mIoU).
4.2 Deployment Details & Parameter Settings
Our implementation is built upon the MMSegmentation [44] codebase. We employ the AdamW optimizer with an initial learning rate of 1e-4, a weight decay of 0.05, epsilon set to 1e-8, and beta parameters of (0.9, 0.999). The total number of training iterations is 40,000, matching REIN, and we adopt a polynomial learning-rate decay schedule that reduces the learning rate to zero over 40,000 iterations with a decay power of 0.9 and no epoch-based warmup. Data augmentation comprises multi-scale resizing, random cropping (with fixed crop size and category-ratio constraint), random horizontal flipping, and photometric distortion. All experiments are conducted on an NVIDIA RTX A6000 GPU with a batch size of 8, taking approximately 14 hours to train and peaking at around 45 GB of GPU memory usage.
4.3 Performance Comparison
Figure 3: Key segmentation examples of existing DGSS methods and Vireo under the Citys. $\rightarrow$ ACDC unseen target domains under Night, Snow, Rain, and Fog conditions.
Table 2: Performance comparison between our Vireo and existing OVSS methods under Citys. $\rightarrow$ DELIVER + Ade150 + Ade847 generalization setting. Top three results are highlighted as best, second, and third, respectively. (%)
Method Proc. & Year Sun-DELIVER Night-DELIVER Cloud-DELIVER Rain-DELIVER Fog-DELIVER Ade150 Ade847
Train on Cityscapes
OVSS Method:
FC-CLIP [29] NeurIPS2023 16.93 14.93 17.50 16.59 17.26 16.12 6.29
EBSeg [30] CVPR2024 26.41 15.50 22.62 20.35 22.00 12.75 3.75
CAT-Seg [10] CVPR2024 28.21 20.56 26.22 26.53 24.80 20.19 6.95
SED [3] CVPR2024 27.14 22.79 24.40 25.18 25.25 18.86 5.45
OV-DGSS Method:
Vireo (Ours) 35.73 27.51 32.34 31.80 32.72 21.37 7.31
Train on GTA
OVSS Method:
FC-CLIP [29] NeurIPS2023 22.24 18.58 18.50 16.59 19.12 15.47 5.73
EBSeg [30] CVPR2024 32.32 20.05 26.19 26.19 28.69 11.87 4.19
CAT-Seg [10] CVPR2024 28.59 23.49 27.31 27.94 27.66 20.45 7.18
SED [3] CVPR2024 26.56 21.18 24.95 24.58 26.17 19.57 6.80
OV-DGSS Method:
Vireo (Ours) 38.49 29.89 33.89 33.46 35.80 21.23 7.68
Domain Generalization Ability: Table 1 summarizes the evaluation results of various state-of-the-art open-vocabulary semantic segmentation (OVSS) and domain-generalized semantic segmentation (DGSS) methods under two cross-domain settings (Cityscapes $\rightarrow$ ACDC, BDD100k, Mapillary, GTA5 and GTA5 $\rightarrow$ Cityscapes, BDD100k, Mapillary). The results demonstrate that our approach achieves outstanding performance across all target datasets, significantly outperforming other OVSS/DGSS methods. Furthermore, the visualizations in Figure 3 shows that Vireo delivers satisfactory predictions in both extreme weather conditions and scenes with dense pedestrian and vehicular traffic.
Open-Vocabulary Capability: Table 2 presents a comparison between Vireo and other open-vocabulary semantic segmentation (OVSS) methods under the Cityscapes → DELIVER (sun, rain, night, cloud, fog), ADE150, and ADE847 configurations. The results indicate that conventional OVSS approaches suffer a sharp performance drop in extreme scenarios (e.g., night), whereas our model—enhanced by depth-based geometric features—maintains robust performance, outperforming the strongest OVSS baseline by at least 5%. Furthermore, as shown in Figure 4, the coexistence of new open classes and extreme weather conditions in the DELIVER dataset leads OVSS approaches to exhibit significant false positives and false negatives.
Table 3 compares Vireo trained on Cityscapes with other open-vocabulary semantic-segmentation methods on the DELIVER dataset across five weather conditions for both Seen and Unseen categories. Vireo achieves the highest mIoU for both groups: it outperforms the second-best method by roughly 7–10 percentage points on Seen classes and by 2–3 points on Unseen ones. Although all methods register lower mIoU on Unseen categories—underscoring the difficulty of open-vocabulary segmentation—Vireo substantially reduces this gap, confirming that its depth-geometry guidance and cross-domain alignment improve recognition of new classes. Overall, Vireo remains consistently superior across weather scenarios and category types, demonstrating stronger out-of-domain generalization.
Table 3: Comparison of Seen and Unseen Category mIoU Across Weather Scenarios and Methods.
Class Cloud Fog Night Rain Sun
Ours SED CAT-Seg FC-CLIP EBSeg Ours SED CAT-Seg FC-CLIP EBSeg Ours SED CAT-Seg FC-CLIP EBSeg Ours SED CAT-Seg FC-CLIP EBSeg Ours SED CAT-Seg FC-CLIP EBSeg
Seen 49.14 36.72 43.92 28.07 38.10 49.59 38.28 41.43 26.06 35.76 41.63 34.99 34.89 23.03 26.03 47.15 37.35 41.43 25.59 31.96 55.16 41.68 45.10 26.52 42.63
Unseen 10.96 8.74 3.69 4.05 2.92 12.15 8.40 3.54 4.32 4.48 9.52 7.27 2.30 4.09 2.09 12.27 8.95 5.81 4.60 5.57 11.46 8.12 4.91 4.73 6.14
Mean 33.89 24.40 26.22 17.50 22.62 35.80 25.25 24.80 17.26 22.00 29.89 22.79 20.56 14.93 15.50 33.46 25.18 26.53 16.59 20.35 38.49 27.14 28.21 16.93 26.41
Figure 4: Key segmentation examples of existing OV-DGSS methods comparison on unseen classes and cross domain on Cityscapes → DELIVER.
4.4 Ablation Study
Table 4: Performance comparison of the proposed Vireo against other DGSS methods under the GTA5 $\rightarrow$ Citys.+BDD.+Map. setting.
EVA02 (Large) [45, 46]
Fine-tune Method
Trainable
Params
mIoU Avg.
Citys. BDD. Map.
Full 304.24M 62.1 56.2 64.6 60.9
+AdvStyle[47] 304.24M 63.1 56.4 64.0 61.2
+PASTA[48] 304.24M 61.8 57.1 63.6 60.8
+GTR-LTR[49] 304.24M 59.8 57.4 63.2 60.1
Freeze   0.00M 56.5 53.6 58.6 56.2
+AdvStyle[47]   0.00M 51.4 51.6 56.5 53.2
+PASTA[48]   0.00M 57.8 52.3 58.5 56.2
+GTR-LTR[49]   0.00M 52.5 52.8 57.1 54.1
+LoRA[50]   1.18M 55.5 52.7 58.3 55.5
+AdaptFormer[51]   3.17M 63.7 59.9 64.2 62.6
+VPT[52]   3.69M 62.2 57.7 62.5 60.8
+REIN[35]   2.99M 65.3 60.5 64.9 63.6
+FADA[36] 11.65M 66.7 61.9 66.1 64.9
+Vireo (Ours)   3.78M 68.5 62.1 67.4 66.0
DINOv2 (Large) [53]
Fine-tune Method
Trainable
Params
mIoU Avg.
Citys. BDD. Map.
Full 304.20M 63.7 57.4 64.2 61.7
+AdvStyle[47] 304.20M 60.8 58.0 62.5 60.4
+PASTA[48] 304.20M 62.5 57.2 64.7 61.5
+GTR-LTR[49] 304.20M 62.7 57.4 64.5 61.6
Freeze   0.00M 63.3 56.1 63.9 61.1
+AdvStyle[47]   0.00M 61.5 55.1 63.9 60.1
+PASTA[48]   0.00M 62.1 57.2 64.5 61.3
+GTR-LTR[49]   0.00M 60.2 57.7 62.2 60.0
+LoRA[50]   0.79M 65.2 58.3 64.6 62.7
+AdaptFormer[51]   3.17M 64.9 59.0 64.2 62.7
+VPT[52]   3.69M 65.2 59.4 65.5 63.3
+REIN[35]   2.99M 66.4 60.4 66.1 64.3
+FADA[36] 11.65M 68.2 62.0 68.1 66.1
+Vireo (Ours)   3.78M 70.7 62.9 69.6 67.7
Figure 5: Visualizations of attention maps and affinity maps under differnet scenes, where CMPE denotes the Coarse Mask Prior Embedding, GTP represents the GeoText Prompts, respectively.
Figure 6: t-SNE embeddings of the feature space. Left: original source-domain dataset. Right: our Vireo after adaptation on Cityscapes $\rightarrow$ ACDC + BDD 100k + Mapillary. Each point is coloured by its semantic class.
Configurations Snow Night Fog Rain BDD. Map.
REIN [35] 70.6 55.9 79.5 72.5 63.5 74.0
$+$ concat $f^{d}_{i}$ 70.8 56.1 79.4 72.8 63.6 74.2
$+$ Prompt Depth Anything 70.4 55.5 79.7 72.0 63.8 73.9
$+$ Depth Anything V2 71.5 56.7 80.5 73.3 64.4 74.5
$+$ DA + AO 72.2 57.4 80.9 74.2 65.1 75.0
$+$ DOV-VEH 70.9 56.2 79.8 72.8 63.7 74.2
$+$ CMPE 71.6 56.9 80.2 73.4 64.1 74.6
$+$ GeoText Prompts 74.0 58.4 81.1 74.8 65.3 75.3
Vireo 76.2 60.6 82.3 76.3 66.7 76.0
Table 5: Ablation studies on component configurations of the proposed Vireo under the Citys. $\rightarrow$ ACDC with Snow, Night, Fog, Rain + BDD. + Map. generalization setting. Top three results are highlighted as best, second, and third (%).
CLIP (ViT-L) [54]
Fine-tune Trainable Params mIoU Avg.
Citys BDD Map
Full 304.15M 51.3 47.6 54.3 51.1
Freeze   0.00M 53.7 48.7 55.0 52.4
REIN   2.99M 57.1 54.7 60.5 57.4
FADA 11.65M 58.7 55.8 62.1 58.9
Vireo (Ours)   3.78M 60.5 57.5 64.1 60.7
SAM (Huge) [55]
Fine-tune Trainable Params mIoU Avg.
Citys BDD Map
Full 632.18M 57.6 51.7 61.5 56.9
Freeze   0.00M 57.0 47.1 58.4 54.2
REIN   4.51M 59.6 52.0 62.1 57.9
FADA 16.59M 61.0 53.2 63.4 60.0
Vireo (Ours)   5.30M 64.5 59.0 66.0 63.2
EVA02 (Large) [45, 46]
Fine-tune Trainable Params mIoU Avg.
Citys BDD Map
Full 304.24M 62.1 56.2 64.6 60.9
Freeze   0.00M 56.5 53.6 58.6 56.2
REIN   2.99M 65.3 60.5 64.9 63.6
FADA 11.65M 66.7 61.9 66.1 64.9
Vireo (Ours)   3.78M 68.5 62.1 67.4 66.0
DINOv2 (Large) [53]
Fine-tune Trainable Params mIoU Avg.
Citys BDD Map
Full 304.20M 63.7 57.4 64.2 61.7
Freeze   0.00M 63.3 56.1 63.9 61.1
REIN   2.99M 66.4 60.4 66.1 64.3
FADA 11.65M 68.2 62.0 68.1 66.1
Vireo (Ours)   3.78M 70.7 62.9 69.6 67.7
Table 6: Performance comparison of Vireo across multiple VFMs under the GTA5 $\rightarrow$ Citys. + BDD. + Map. setting. Trainable parameters in backbone. Top three results are best, second, third. (%)
[…]
Figure 7: Comparison of Train Loss for Baseline and CMPE Models.
Table 6 reports the parameter overhead and mIoU of Vireo and several lightweight fine-tuning schemes on four backbones (CLIP-L, SAM-H, EVA02-L, and DINOv2-L). Compared with the prompt-based REIN baseline, Vireo adds only about 0.79 M extra parameters yet achieves the highest average mIoU on every backbone: the gain is most pronounced on the parameter-constrained CLIP-L, and Vireo still surpasses heavier adapters such as FADA by 1–2 mIoU on the larger EVA02-L and DINOv2-L models, underscoring its parameter efficiency and scalability across backbones.
In Table 5, using Depth Anything V2 alone yields an approximate $0.6\%$ mIoU gain across all six scenarios, and adding depth augmentation with attention optimization (DA + AO) delivers a further $\approx 1\%$ improvement. The DOV-VEH and CMPE modules each add $0.5\%$$0.8\%$, validating mask vectors and dense gradient embedding. GeoText Prompts on its own provides a substantial $\approx 4.4\%$ boost, underscoring the complementary benefits of fusing semantic and geometric cues.
Pronounced Attention Focus. Figure 5 further confirm that GeoText Prompts steer the model toward geometry-sensitive regions, while CMPE strengthens gradients on foreground masks; compared with baselines, Vireo’s focus is tighter on scene structure and semantic boundaries, explaining its consistent advantage in cross-domain and open-vocabulary settings.
t-SNE Visualization of DGSS Features. The feature distributions of the original dataset and our method are visualized in Figure 6, revealing the superiority of our learned features in forming well-separated semantic clusters. This demonstrates the effectiveness of our domain-generalized visual-textual alignment in structuring the open-vocabulary semantic space.
5 Conclusion
This study introduces Vireo, the first single-stage framework that unifies open-vocabulary recognition and domain-generalised semantic segmentation. By integrating frozen visual foundation models, depth-aware geometry and three core modules—GeoText Prompts, Coarse Mask Prior Embedding and the DOV Vector Embedding Head—it converges faster, concentrates attention on scene structure and outperforms state-of-the-art methods across multiple benchmarks, demonstrating the power of combining textual cues with geometric priors for robust pixel-level perception.
Limitations Vireo assumes a reliable RGB camera, but in rare cases (e.g., occlusion, glare, hardware failure), the stream can be lost, impairing perception. Future work will explore multi-source setups—automatically switching to lidar, radar, or event-based cameras when RGB fails—to keep segmentation robust in all conditions.
Appendix / Supplemental Material
Appendix A Open Source Dataset and Experiments
To advance research in real-world open-vocabulary and domain-generalized semantic segmentation, we introduce AustinScapes, a new large-scale street-view dataset collected from the city of Austin, Texas, USA. The dataset comprises 100,000 high-resolution street-level images, each accompanied by pixel-level semantic annotations. The annotation schema follows the Cityscapes 19-class labels, ensuring compatibility with existing benchmarks.
AustinScapes captures a wide variety of urban scenes under diverse lighting, architectural styles, and traffic conditions, making it well-suited for evaluating both fine-grained recognition and robustness across real-world complexities. We benchmark our method and previous state-of-the-art approaches on AustinScapes to assess generalization performance. Figure 8 shows the visualization results of DGSS and OVSS methods on the AustinScapes dataset. Experimental results show that our framework maintains superior segmentation quality, particularly in challenging object categories and under domain shifts, further validating its effectiveness on unseen urban domains.
Figure 8: Key zero-shot segmentation examples from existing DGSS and OVSS methods, and our Vireo, on the AustinScapes dataset.
Appendix B Related Works
This section provides a more comprehensive discussion of the research areas that are highly relevant to our work: Open-Vocabulary Semantic Segmentation (OVSS) and Domain-Generalized Semantic Segmentation (DGSS). By examining these areas, we aim to highlight our contributions in bridging the gap between DGSS and OVSS.
B.1 Open-Vocabulary Semantic Segmentation
Existing OVSS methods [24, 10, 30, 3] primarily leverage the rich semantic knowledge embedded in pre-trained vision-language models (VLMs) such as CLIP [54]. Early approaches often adopt two-stage pipelines. For example, ZSSeg [56] classifies class-agnostic masks using CLIP, and MaskCLIP [23] adapts CLIP for instance mask proposal classification. Towards end-to-end solutions, LSeg [22] generates pixel-level embeddings aligned with text features, and OVSeg [1] further enhances this by learning discriminative pixel embeddings using a class-agnostic decoder. OpenSeg [57] presents a unified, language-driven framework for open-vocabulary segmentation.
To reduce dependency on region proposals, FreeSeg [2] introduces a grouping-based strategy to cluster pixels. Generalized decoders such as X-Decoder [58] enable language-conditioned pixel-level predictions. GroupViT [59] explores hierarchical region grouping with textual concepts. SegCLIP [11], a task-specific adaptation, enhances spatial understanding from CLIP for dense prediction tasks. More recently, OpenSeeD [60] aims to unify multiple segmentation tasks under the open-vocabulary setting via a conditioned decoder, and Cat-seg [10] improves multimodal matching by jointly modeling image and text in the cost function. Additionally, [61] addresses evaluation issues by proposing a mask-matching-based framework to handle semantic ambiguity.
Despite these advances, challenges remain in transferring coarse-grained VLM knowledge to pixel-level predictions and in resolving fine-grained linguistic ambiguities. Our work builds upon this line of research by incorporating depth cues to enhance language-guided segmentation, particularly in disambiguating objects within complex scenes.
B.2 Domain-Generalized Semantic Segmentation
Previous methods primarily focus on learning feature representations that are both discriminative for semantic segmentation and invariant to domain-specific variations. A foundational strategy in DGSS is introduced by DANN [27], which employs adversarial training with a domain classifier. Li et al. [62] investigate generative approaches using GANs to model the joint distribution of images and labels, while RobustNet [28] mitigates domain-specific biases through instance-selective whitening. SAN-SAW [13] introduces semantic-aware normalization using SAN and SAW modules, and BlindNet [63] enhances robustness by disentangling style and content representations.
More recently, pre-trained VFMs have been effectively leveraged in DGSS frameworks [8, 6, 64, 5]. These methods utilize frozen VFM features as universal backbones and optimize learnable tokens to improve cross-domain performance.
Building upon these advancements, our work posits that integrating the semantic generality of language with the geometric stability of depth provides a compelling approach to achieving robust domain-generalized open-vocabulary segmentation.
Appendix C Expend Experiments and Analysis
In this section, we present per-class method comparisons, model details, and visual results that are excluded from the main manuscript owing to page constraints.
C.1 More Data Analysis and Model Details
From Table 7, our approach achieves a mIoU of 66.73%, significantly outperforming other OVSS approaches. Compared to the second-best SED (48.16% mIoU), this represents an improvement of 18.57 percentage points, demonstrating a substantial enhancement in cross-domain generalization. Moreover, our approach attains the best performance in 18 categories.
For small targets prone to missed detections under extreme conditions (such as traffic lights and pedestrians), our method achieves 62.31% and 72.94% (17.34% and 12.88% higher than SED), respectively. The results reflect the robustness of our method in complex scenarios.
Table 7: Performance comparison across OVSS methods on per-class IoU and mIoU (City.→BDD.). Top three results are highlighted as best, second, and third, respectively(%).
Method road side. build. wall fence pole light sign vege terr. sky pers. rider car truck bus train moto. bicy. mIoU
EBseg 89.72 42.47 81.24 25.11 31.24 45.95 40.14 44.32 80.57 17.67 93.20 50.52 24.96 83.97 35.60 62.83 0.00 39.60 40.02 40.59
FC-CLIP 93.63 57.06 79.53 18.63 38.78 48.86 24.17 48.45 83.89 45.27 83.26 69.16 55.44 88.55 44.13 75.35 0.00 66.74 60.34 51.41
CAT-Seg 92.08 50.94 79.41 29.08 38.85 13.88 25.88 28.88 79.26 44.20 91.11 45.51 30.07 82.74 40.65 64.85 0.00 43.23 36.18 44.28
SED 93.78 60.35 82.90 30.07 44.48 39.26 44.97 46.22 83.40 45.55 92.53 60.06 29.23 86.92 43.64 42.94 0.31 43.88 42.12 48.16
Ours 95.40 68.08 88.29 34.81 46.11 60.95 62.31 59.80 87.26 51.10 95.27 72.94 62.63 90.93 53.82 83.08 22.35 74.40 58.37 66.73
Table 8: Comparison of per-class IoU across OVSS methods on the DELIVER dataset. Top three results are highlighted as best, second, and third, respectively(%).
Class Cloud Fog Night Rain Sun
Ours SED CAT-Seg FC-CLIP EBseg Ours SED CAT-Seg FC-CLIP EBseg Ours SED CAT-Seg FC-CLIP EBseg Ours SED CAT-Seg FC-CLIP EBseg Ours SED CAT-Seg FC-CLIP EBseg
Building 73.71 81.43 81.63 53.96 81.64 52.58 64.43 59.10 50.79 62.79 56.24 77.20 74.99 58.13 75.90 57.34 72.25 79.73 60.24 71.22 85.55 83.17 81.47 49.24 84.39
Fence 39.24 18.22 25.89 28.31 29.54 37.52 14.10 19.50 25.53 21.56 21.20 7.07 10.65 12.99 8.94 38.06 11.77 16.88 24.36 20.86 37.10 10.67 23.77 22.65 28.26
Other 0.00 0.00 0.00 0.11 0.04 0.00 0.00 0.00 0.02 0.06 0.00 0.00 0.00 0.00 0.13 0.00 0.00 0.00 0.00 0.17 0.00 0.00 0.00 1.88 0.03
Pedestrian 68.86 39.43 21.84 4.21 16.64 73.61 36.72 14.02 2.96 36.21 56.44 31.99 9.00 8.78 5.93 66.21 33.65 19.85 3.28 17.14 64.61 30.83 13.25 3.37 23.59
Pole 32.17 19.50 15.62 6.74 31.35 31.91 21.96 17.90 5.82 35.65 24.20 20.46 14.19 6.51 33.20 32.85 26.89 20.69 6.69 34.53 37.35 22.33 19.39 4.20 34.08
RoadLine 0.00 5.87 2.30 4.46 2.95 0.00 5.44 2.28 3.94 1.30 0.00 5.09 1.94 3.59 2.68 0.00 5.63 2.11 4.01 2.38 0.00 4.80 2.09 3.96 3.02
Road 91.70 29.72 89.55 0.00 65.80 91.72 24.25 89.15 0.01 80.96 90.78 26.22 86.28 0.00 64.23 91.65 32.11 88.48 0.01 67.85 91.94 22.69 88.78 0.00 63.06
SideWalk 71.50 57.22 57.36 59.27 59.04 66.91 53.90 55.25 47.48 43.27 69.18 56.95 58.65 54.13 38.89 69.50 56.40 56.95 52.37 45.44 71.16 48.96 55.14 44.70 27.36
Vegetation 69.36 61.11 62.12 29.30 39.67 68.58 58.86 58.96 33.79 35.20 38.46 40.39 31.46 24.22 20.33 66.69 62.29 62.37 28.17 31.89 68.85 58.12 59.30 54.17 68.62
Cars 63.84 60.20 57.84 66.49 65.73 68.99 60.90 58.78 72.55 68.84 77.26 66.19 60.66 74.09 67.51 64.80 59.94 53.38 62.80 60.64 70.38 65.32 59.65 66.64 64.05
Wall 3.46 0.08 2.05 2.67 4.14 0.30 0.00 0.00 0.08 0.43 2.11 1.46 0.00 1.47 1.53 0.47 0.41 0.01 0.49 0.23 4.07 5.95 0.00 0.29 3.80
TrafficSign 22.00 26.99 25.02 30.59 23.55 32.28 36.64 24.98 10.11 25.88 20.83 35.63 33.87 12.18 26.76 13.01 28.84 26.61 14.26 20.98 32.60 41.36 42.29 29.98 32.87
Sky 86.90 92.22 94.09 26.28 59.90 67.51 82.47 82.87 18.12 41.63 43.22 79.51 56.92 21.98 7.81 65.98 77.80 92.65 11.00 13.61 96.37 95.20 95.22 23.70 97.22
Ground 0.00 0.01 0.00 0.00 0.01 0.00 0.00 0.00 0.00 0.00 0.00 0.05 0.00 0.00 0.00 0.00 0.00 0.07 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Bridge 0.00 0.00 0.15 0.00 0.00 0.00 0.00 0.11 0.00 0.00 0.00 0.00 0.04 0.13 0.00 0.00 0.00 0.01 0.03 0.00 0.00 0.00 0.29 0.00 0.00
RailTrack 0.06 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.21 0.00 0.00 0.00 0.00 0.06 0.00 0.00 0.00 0.25 0.00 0.00 0.00 0.00 0.00
GroundRail 0.00 12.80 0.00 15.07 0.00 0.00 8.32 0.03 11.33 0.00 0.00 11.20 0.00 6.31 0.00 0.00 6.93 0.00 8.85 0.00 0.00 6.72 0.00 11.54 0.00
TrafficLight 57.73 25.71 23.39 24.94 13.05 67.39 41.82 29.52 10.24 21.85 51.48 30.92 21.47 2.01 4.45 58.03 23.86 22.72 4.60 10.50 61.49 39.72 25.74 3.17 17.78
Static 0.00 0.07 0.00 0.12 0.00 0.00 0.01 0.01 0.00 0.22 0.00 0.01 0.00 0.00 0.00 0.00 0.00 0.01 0.00 0.00 0.00 0.00 0.04 0.02 0.00
Dynamic 0.00 0.00 0.00 0.01 0.36 0.00 0.00 0.00 0.00 0.06 0.00 0.00 0.00 0.00 0.53 0.00 0.00 0.00 0.01 0.18 0.00 0.00 0.00 0.00 0.78
Water 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.68 0.00 0.00 0.00 0.38 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Terrain 48.31 32.42 38.57 15.22 4.35 45.35 27.11 35.20 17.08 2.42 37.71 13.06 28.41 4.41 0.45 47.53 25.24 33.96 17.23 9.46 43.44 25.18 31.03 15.49 5.16
TwoWheeler 51.66 37.99 16.31 20.55 12.10 60.03 41.90 22.45 29.29 11.42 48.04 31.61 14.29 25.50 13.74 68.66 52.24 41.47 34.45 41.11 60.42 47.01 38.37 31.28 40.16
Bus 0.00 0.01 0.00 3.76 0.00 0.00 6.11 0.00 4.45 0.00 0.00 0.93 0.00 0.00 0.00 0.00 0.00 0.00 23.64 12.30 0.00 0.00 0.00 1.73 0.05
Truck 28.00 9.22 41.70 45.42 55.62 53.19 43.36 49.80 68.80 60.14 50.18 33.77 10.99 50.29 14.40 54.24 45.11 45.65 52.43 47.98 67.95 64.92 69.59 55.30 66.04
Mean 33.89 24.40 26.22 17.50 22.62 35.80 25.25 24.80 17.26 22.00 29.89 22.79 20.56 14.93 15.50 33.46 25.18 26.53 16.59 20.35 38.49 27.14 28.21 16.93 26.41
Table 9: Efficiency and Performance Comparison
Method Iter. Train (4 BS) Mem. FPS Infer. Perform.
SED 40k 6h59min 20.8G 1.73 33min 57.0%
CAT-Seg 40k 3h3min 9.7G 1.66 35min 48.9%
FC-CLIP 40k 5h34min 18.7G 1.45 40min 58.7%
EBSeg 40k 6h34min 22.1G 0.62 94min 49.7%
REIN 40k 12h6min 13.8G 1.88 31min 69.4%
Vireo 40k 15h2min 18.4G 1.49 39min 73.0%
As shown in Table 8, on the DELIVER dataset, our method achieves the highest mIoU across all five extreme weather conditions, outperforming the runner-up by an average of 8–11 percentage points. Under Sun and Cloud conditions, large classes achieve over 80% IoU. In more adverse conditions, our method still maintains IoUs above 30%. Even in the most challenging Night scenario, we outperform others by a margin of 7.10%. For fine-grained categories such as TwoWheeler, Pedestrian, TrafficSign, and TrafficLight, our method consistently outperforms competitors, demonstrating strong recognition of low-contrast and small-scale objects. Although rare categories exhibit lower absolute IoUs, we still rank among the top in most of these cases. Overall, our approach demonstrates superior performance across both frequent and challenging classes under diverse weather conditions.
Table 9 compares training time, memory usage, FPS, inference time, and final mIoU (averaged across ACDC, BDD100K, and Mapillary) for various OVSS methods and the baseline, all evaluated under a unified 40k-iteration schedule with a batch size of 8. While Vireo’s training time (15 hours 2 minutes) and inference time (43 minutes) are not the shortest, it achieves the highest mIoU of 73.0%, utilizing 18.4 GB of memory at 1.36 FPS, substantially outperforming all other approaches.
Importantly, we follow the previous configuration of intermediate DINOv2 feature layers ($\{\hat{f}^{V}_{l_{i}}\}_{i=1}^{4},l_{i}\in\{8,12,16,24\}$), which has been demonstrated in prior work to offer an optimal trade-off between accuracy and efficiency. This ensures a fair and consistent comparison across all methods in our experiments.
C.2 More Visualization Results
Figure 9 presents a visual comparison of segmentation results between DGSS methods and ours on the Cityscapes $\rightarrow$ ACDC target domain. The first row shows the original outputs, while the second and third rows display zoomed-in crops for closer inspection. In the original images, other models clearly miss regions on the left sidewalk. In the second-row crops, our method yields much finer and more complete boundaries around small objects such as traffic lights and signs, detecting nearly every instance. In the third-row crops, even distant vehicles and signals are accurately segmented. Moreover, Vireo maintains continuous, clean segmentation of roads and sidewalks in low-contrast conditions (nighttime, rain, snow), without breaks or noise artifacts. It preserves highly consistent segmentation across large background areas like sky, buildings, and vegetation, suppressing “speckle” noise. It precisely captures very fine structures such as curbs and road markings. And it produces coherent, complete contours for dynamic objects (pedestrians, cyclists, vehicles). Whether in clear daytime or foggy, rainy nighttime scenarios, Vireo demonstrates superior cross-scenario robustness.
Figure 9: Key segmentation examples of existing DGSS methods and Vireo under the Citys. $\rightarrow$ ACDC unseen target domains under Night, Snow, Rain, and Fog conditions.
Figure 10 presents a visual comparison of segmentation results between DGSS methods and ours on the GTA5 $\rightarrow$ Cityscapes + BDD + Mapillary target domains. Across these varying scene distributions, RobustNet, HGFormer, and REIN often misclassify or omit road markings, crosswalks, clusters of pedestrians, and parked vehicles. In contrast, our Vireo consistently preserves the complete contours of dynamic objects and slender structures (e.g., lampposts, traffic signals). It segments curbs and lane lines continuously without breaks and delineates building–sky boundaries more sharply. It renders vegetation and other background regions with uniform, noise-free color blocks. Even for distant groups of pedestrians or in highly cluttered street scenes, Vireo accurately separates adjacent classes. This further demonstrates its superior generalization and robustness to unseen domains.
Figure 10: Key segmentation examples of existing DGSS methods and Vireo under the GTA5→City., BDD. and Map. unseen target domains.
Figure 11 presents a comparison of segmentation results between various OVSS methods and Vireo on the Cityscapes $\rightarrow$ ACDC target domain. The first column shows the original input images, followed by CAT-Seg, EBSeg, FC-CLIP, SED, Vireo, and finally the ground truth. As observed, CAT-Seg and EBSeg often yield fragmented road and sidewalk segmentations with background noise under extreme conditions such as rain, snow, and nighttime. FC-CLIP partially recovers large objects but still fails to detect small targets like traffic lights and thin poles. SED shows some improvement in road continuity but struggles with distant small objects and background consistency. In contrast, Vireo consistently produces continuous, precise segmentations of roads, lane markings, pedestrians, vehicles, and traffic signals across all weather and lighting conditions. It maintains low-noise, highly uniform background regions. Its outputs most closely match the ground truth, further demonstrating Vireo’s superior performance and robustness in open-vocabulary, cross-domain scenarios.
Figure 11: Key segmentation examples of existing OVSS methods and Vireo under the Citys. $\rightarrow$ ACDC unseen target domains under Night, Snow, Rain, and Fog conditions.
Figure 12 compares OVSS methods and our Vireo on the Cityscapes $\rightarrow$ DELIVERtransfer setting, which includes sunny, rainy, cloudy, foggy, and nighttime conditions. FC-CLIP preserves mask edges but suffers from widespread class errors. CAT-Seg and EBSeg yield fragmented road and sidewalk masks with spurious background artifacts under adverse conditions. SED improves overall continuity but still missegments slender structures like poles and signs in low-contrast scenarios. In contrast, Vireo consistently produces clear, coherent masks for zebra crossings, lane markings, sidewalks, pedestrians, vehicles, and poles. It effectively suppresses noise in sky, building, and vegetation regions. The outputs most closely match the ground truth, demonstrating superior robustness and generalization in extreme, unseen domains.
As shown in Figure 13, adding CMPE and GTP transforms the attention from a diffuse pattern into a concentrated focus on critical regions such as roads, vehicles, and pedestrians. In the Affinity Map, bright clusters reveal the model’s ability to capture long-range dependencies within the same object, while dark boundaries emphasize clear separation between different semantic regions. Together, these results confirm our method’s robust handling of semantic coherence and pixel-level affinities across diverse weather and lighting conditions.
Figure 12: Key segmentation examples of existing OVSS methods and Vireo under the Citys. $\rightarrow$ DELIVER unseen target domains under Cloud, Fog, Night, Rain, and Sun conditions.
Figure 13: Visualizations of attention maps and affinity maps under different scenes, where CMPE denotes the Coarse Mask Prior Embedding and GTP represents the GeoText Prompts.